A zoo has no lions and ostriches. Pet Sitting guy spoke 15 heads and 50 legs. How many lions?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)


A zoo has no lions and ostriches. Pet Sitting guy spoke 15 heads and 50 legs. How many lions?
-------------
None, you just said no lions.

Answer by stanbon(75887)

A zoo has lions and ostriches. Pet Sitting guy sees 15 heads and 50 legs.
How many lions are there?
-------
Head Equation: L + O = 15
Legs Equation:4L +2O = 50
-------------------------------
Modify for eliminatiion:
4L + 4O = 4*15
4L + 2O = 50
-----
Subtract and solve for "O":
2O = 10
O = 5 (# of ostriches)
-----
Solve for "L"
L + O = 15
L + 5 = 15
L = 10 (# of lions)
------------------------
cheers,
Stan H.
========================